The number of visitors to a website in the first week is 50. The number of visitors each week is double the number of visitors the previous week. What is the total number of visitors to the website in the first 8 wk?

Answer by ikleyn(52788)

.
Sum of the 8 first terms of the geometric progression with the first term 50 and the common ratio of 2:


  50 + 5*2 + 50*2^2 + . . . + 50*2^7 = 


= 50*(1 + 2 + 2^2 + . . . + 2^7) = 50*(2^8-1) = 50*(256-1) = 50*255 = 12750.

------------------
On geometric progressions, see introductory lessons in this site
    - Geometric progressions
    - The proofs of the formulas for geometric progressions
    - Problems on geometric progressions
    - Word problems on geometric progressions
    - Solved problems on geometric progressions
    - Fresh, sweet and crispy problem on arithmetic and geometric progressions

Also,  you have this free of charge online textbook in ALGEBRA-II in this site
    - ALGEBRA-II - YOUR ONLINE TEXTBOOK.

The referred lessons are the part of this online textbook under the topic "Geometric progressions".


Save the link to this textbook together with its description

Free of charge online textbook in ALGEBRA-II
https://www.algebra.com/algebra/homework/complex/ALGEBRA-II-YOUR-ONLINE-TEXTBOOK.lesson

into your archive and use when it is needed.